# Extending MGCAMB tests of gravity to nonlinear scales

### v1

@@ Abstract
Modified Growth with CAMB (MGCAMB) is a patch for the Einstein-Boltzmann solver CAMB for cosmological tests of gravity. Until now, MGCAMB was limited to scales well-described by linear perturbation theory. In this work, we extend the framework with a phenomenological model that can capture nonlinear corrections in a broad range of modified gravity theories. The extension employs the publicly available halo model reaction code ReACT, developed for modeling the nonlinear corrections to cosmological observables in extensions of the $\Lambda$CDM model. The nonlinear extension makes it possible to use a wider range of data from large scale structure surveys, without applying a linear scale cut. We demonstrate that, with the 3$\times$2pt Dark Energy Survey data, we achieve a stronger constraint on the linear phenomenological functions $\mu$ and $\Sigma$, after marginalizing over the additional nonlinear parameter $p_{1}$, compared to the case without the nonlinear extension and using a linear cut. The new version of MGCAMB is now forked with CAMB on GitHub allowing for compatibility with future upgrades.
@@ II.1 The halo model reaction method
$\mathcal{R}(k,z)\equiv\frac{P_{\text{NL}}(k,z)}{P_{\mathrm{L}}^{\mathrm{pseudo}}(k,z)+P_{1h}^{\mathrm{pseudo}}(k,z)}\,,$ (6)
Here, $P_{\text{NL}}(k,z)$ is the nonlinear power spectrum for the target cosmology, whereas the denominator contains the nonlinear ‘pseudo’ spectrum, which is defined as the $\Lambda$CDM power spectrum whose linear clustering matches the modified cosmology at the target redshift $z$, i.e.:
@@
In Eq. (6) we have implicitly assumed the halo model to construct the pseudo power spectrum, where we have separated the 2-halo, or linear, term in the power spectrum $P_{\mathrm{L}}^{\mathrm{pseudo}}(k,z)$, and the 1-halo term, $P_{1h}^{\mathrm{pseudo}}(k,z)$. The reaction function represents all the corrections to the pseudo power spectrum coming from nonlinear beyond-$\Lambda$CDM physics.
Based on the approach in Refs. Cataneo et al. (2020); Bose et al. (2020, 2022), the reaction function can be computed as:
$\mathcal{R}(k)=$
$\frac{\left(1-f_{\nu}\right)^{2}P_{\mathrm{HM}}^{(\mathrm{cb})}(k)+2f_{\nu}\left(1-f_{\nu}\right)P_{\mathrm{HM}}^{(\mathrm{cb}v)}(k)+f_{\nu}^{2}P_{\mathrm{L}}^{(\nu)}(k)}{P_{\mathrm{HM}}^{\mathrm{pseudo}}(k)}\,,$ (8)
where ‘HM’ represents the halo model, ‘cb’ denotes cold dark matter and baryons, ‘$\nu$’ denotes massive neutrinos, $f_{\nu}=\Omega_{\nu,0}/\Omega_{\rm m,0}$ is the energy density fraction of massive neutrinos to the total matter density today, and we have omitted the $z$ dependence for compactness of notation. In Eq. (8), the total matter power spectrum is constructed from the weighted sum of individual components in the presence of massive neutrinos Agarwal and Feldman (2010). $P_{\mathrm{HM}}^{(\mathrm{cb}\nu)}(k)$ stands for the cross power spectrum between neutrinos and the other two matter components, defined as
$P_{\mathrm{HM}}^{(\mathrm{cb}\nu)}(k)\approx\sqrt{P_{\mathrm{HM}}^{(\mathrm{cb})}(k)P_{\mathrm{L}}^{(\nu)}(k)}\,,$ (9)
where $P_{\mathrm{HM}}^{(\mathrm{cb})}$ is given by Cooray and Sheth (2002); Bose et al. (2022)
$P_{\mathrm{HM}}^{(\mathrm{cb})}(k)=\left[(1-\mathcal{E})e^{-k/k_{\star}}+\mathcal{E}\right]P_{\mathrm{L}}^{(\mathrm{cb})}(k)+P_{1h}^{(\mathrm{cb})}(k)\,.$ (10)
In these equations, we have defined
$\mathcal{E}(z)=$ $\lim_{k\rightarrow 0}\frac{\left(1-f_{\nu}\right)^{2}P_{1h}^{(\mathrm{cb})}(k,z)}{P_{1h}^{\mathrm{pseudo}}(k,z)}\,,$ (11)
$k_{\star}(z)=$ $-\bar{k}\left(\ln\left[\frac{A_{1}(\bar{k},z)\pm A_{2}(\bar{k},z)}{\left(1-f_{\nu}\right)^{2}P_{\mathrm{L}}^{(\mathrm{cb})}(\bar{k},z)\big{(}1-\mathcal{E}(z)\big{)}}\right]\right)^{-1}\,,$ (12)
with $A_{1}(k,z)$ and $A_{2}(k,z)$ expressed as
$A_{1}(k,z)=$ $f_{\nu}^{2}P_{\mathrm{L}}^{(\nu)}(k,z)+P_{\mathrm{HM}}^{\mathrm{pseudo}}(k,z)\mathcal{R}_{\mathrm{SPT}}(k,z)$
$-\left(1-f_{\nu}\right)^{2}\left[\mathcal{E}(z)P_{\mathrm{L}}^{(\mathrm{cb})}(k,z)+P_{1h}^{(\mathrm{cb})}(k,z)\right]\,,$ (13)
$A_{2}(k,z)=$ $2\sqrt{f_{\nu}^{2}P_{\mathrm{HM}}^{\mathrm{pseudo}}(k,z)P_{\mathrm{L}}^{(\nu)}(k,z)\mathcal{R}_{\mathrm{SPT}}(k,z)}\,,$ (14)
respectively, and the default scale where we calculate $k_{\star}$ is set to $\bar{k}=0.06$ $h$ Mpc-1.
$\mathcal{R}_{\mathrm{SPT}}(k,z)$ is the 1-loop standard perturbation theory (SPT) prediction for the reaction function, which characterizes the quasi-nonlinear scales well, and is used to calibrate the halo model reaction to better predict this transitory regime in the presence of modifications to gravity. Typically SPT has a good accuracy in the range $0.10.01<k~{}\mathrm{Mpc}\,h^{-1}\,<$ 0.1, whereas the nonlinear regime is typically considered to be given by $k>0.1~{}h\,\mathrm{Mpc}^{-1}$ for late time structure formation. This latter regime can be more accurately predicted by $P_{1h}^{(\mathrm{cb})}(k)$.
@@ II.2 The nonlinear phenomenological parameterization
There are four additional parameters in Eqs. 17-19, characterising the nonlinear regime, $p_{1}$ $p_{4}$, representing MG phenomenological effects. The first one, $p_{1}$, parameterizes the strength of screening, $p_{2}$ and $p_{3}$ quantify the mass and environmental dependencies, respectively, whereas $p_{4}$ sets the scale of Yukawa suppression Yukawa (1935). The QNL correction, described by $S(k,a)$, was shown in Bose et al. (2022) to have a negligible effect for scale-independent MG theories, while it can have up to a 2% effect on the power spectrum for the Hu-Sawicki $f(R)$ model. We will ignore the QNL correction in the current implementation.
@@ IV Demonstrations
Parameters Flat Prior Fiducial
@@
First, we perform a test on mock data, generating synthetic DES-like galaxy-galaxy, galaxy-lensing and lensing-lensing (3$\times$2pt) correlations for a fiducial model with $\mu_{0}=0.4,\Sigma_{0}=0.1$ and $p_{1}=0.5$, and assuming a $\Lambda$CDM background expansion. The values of the other (cosmological) parameters are reported in Tab. 1, whereas the DES likelihood parameters are fixed at the standard values of DES Year-1 Abbott et al. (2019); Krause et al. (2017), to help us isolate the impact of including the nonlinear extension. The mock data are generated using MGCAMB. As our “data” covariance, we adopt the covariance matrix from the DES Year-1 likelihood, rescaled down by a factor of 25, as the actual DES Year-1 covariance would not allow us to constrain the nonlinear parameter $p_{1}$. We then run MCMC chains using MGCobaya, to see if we can recover the input model. To test our implementation, and to assess the added constraints brought in by the nonlinear scales, we run with and without the nonlinear extension. In the latter case, we only vary the linear parameters $\mu_{0}$ and $\Sigma_{0}$, whereas we include $p_{1}$ when the nonlinear extension is applied. In the analyses we run, we employ the priors reported in Tab. 1.
@@
To test the nonlinear extension to MGCAMB, we first check that it does not impact constraints derived from the linear scales alone. To do so, we analyse the DES-like data with and without the nonlinear extension, while applying the aggressive cut. We show the results of this test in Fig. 3 and the first two columns of Tab. 2: the recovered constraints on the parameters $\mu_{0}$ and $\Sigma_{0}$ are consistent with each other and with the fiducial model in both cases, demonstrating that indeed the nonlinear corrections does not impact the linear scales in unwanted ways.
@@
As a second test, we assess the improvement in the constraints when we include the nonlinear scales. For this, we analyse the baseline DES-like synthetic data with the nonlinear extension. We separately analyse the same data with the nonlinear extension excluded, this time without varying $p_{1}$. The joint constraints on $\mu_{0}$, $\Sigma_{0}$ and $p_{1}$ are shown in Fig. 4. It is clear that we obtain better constraints on $\mu_{0}$ and $\Sigma_{0}$ when the nonlinear extension is used, even after marginalizing over $p_{1}$. We report the parameter values we recover in Tab. 2, showing that the $1\sigma$ uncertainty on $\mu_{0}$ is reduced by a factor of $\sim 2$ when the nonlinear extension is used, which agrees with Fig.3. of Ref. Tsedrik et al. (2024).
Figure 4: Comparison of the constraints on $\{\mu_{0},\Sigma_{0}\}$ obtained with and without the nonlinear extension. In the latter case, the aggressive cut is applied. The constraints are derived from synthetic data generated from the fiducial parameters reported in Tab. 1. Constraints are tighter when the nonlinear extension and the baseline data are used.
@@
$\mu_{0}$ 0.003 $\pm$ 0.201 -0.012 $\pm$ 0.216
$\Sigma_{0}$ 0.016 $\pm$ 0.039 0.043 $\pm$ 0.045
Table 3: Parameter inferences on $\mu_{0}$ and $\Sigma_{0}$ ($68\%$ CL) from DES Year-3 baseline data, with and without the nonlinear extension (also see Fig. 5). When excluding the nonlinear extension, we use the same linear cut as the DES collaboration Abbott et al. (2023). PBRS denotes the combination of Planck 2018, BAO, RSD, and SN Ia data, as detailed in Sec. IV. Note the improvement on the constraints when the nonlinear extension is included.
@@
Having tested our setup, we finally apply the extended MGCAMB to current cosmological datasets, deriving constraints on $\mu_{0}$ and $\Sigma_{0}$. For this, we use the DES Year-3 3$\times$2pt data Abbott et al. (2023), Planck 2018 CMB temperature, polarization and lensing power spectra Aghanim et al. (2019), joint measurements of baryon acoustic oscillations (BAO) and redshift-space distortions (RSD) from eBOSS DR16, the SDSS DR7 MGS data Ross et al. (2015), the BAO measurement from 6dF Beutler et al. (2011), and the Pantheon sample of uncalibrated supernovae Scolnic et al. (2018). To assess, again, the impact of including nonlinear scales, we perform two analyses, with and without the nonlinear extension, applying the DES Year-3 linear cut in the latter. We show the recovered posteriors in Fig. 5: it is clear that we obtain stronger constraints on $\mu_{0}$ and $\Sigma_{0}$ with the nonlinear extension. In fact, whilst the posterior on $\mu_{0}$ is largely unaffected, the improvement on the $\Sigma_{0}$ constraint is apparent. We can understand this by noting that DES is primarily a weak lensing survey, which are especially constraining for $\Sigma$. Our parameter estimates are provided in Tab. 3. The constraints we obtain with the DES Year-3 linear cut are consistent with those reported by the DES collaboration Abbott et al. (2023).

### v2

@@ Abstract
Modified Growth with CAMB (MGCAMB) is a patch for the Einstein-Boltzmann solver CAMB for cosmological tests of gravity. Until now, MGCAMB was limited to scales well-described by linear perturbation theory. In this work, we extend the framework with a phenomenological model that can capture nonlinear corrections in a broad range of modified gravity theories. The extension employs the publicly available halo model reaction code ReACT, developed for modeling the nonlinear corrections to cosmological observables in extensions of the $\Lambda$CDM model. The nonlinear extension makes it possible to use a wider range of data from large scale structure surveys, without applying a linear scale cut. We demonstrate that, with the 3$\times$2pt Dark Energy Survey data, we achieve a stronger constraint on the linear phenomenological functions $\mu$ and $\Sigma$, after marginalzing over the additional nonlinear parameter $p_{1}$, compared to the case without the nonlinear extension and using a linear cut. The new version of MGCAMB is now forked with CAMB on GitHub allowing for compatibility with future upgrades.
@@ II.1 The halo model reaction method
$\mathcal{R}(k,z)\equiv\frac{P_{\text{NL}}(k,z)}{P^{\mathrm{pseudo}}(k,z)}\,.$ (6)
Here, $P_{\text{NL}}(k,z)$ is the nonlinear power spectrum for the target cosmology, whereas the denominator contains the ‘pseudo’ spectrum, which is defined as a nonlinear $\Lambda$CDM power spectrum whose linear clustering matches the modified cosmology at the target redshift $z$, i.e.:
@@
We can produce a prediction for Eq. (6) by assuming the halo model to construct the spectra. In this case the spectra are given by a sum of the 2-halo, or linear, term in the power spectrum $P_{\mathrm{L}}(k,z)$, and the 1-halo term, $P_{1h}(k,z)$ and $P_{1h}^{\mathrm{pseudo}}(k,z)$ for MG and pseudo cosmologies respectively. The reaction function represents all the corrections to the pseudo power spectrum coming from nonlinear beyond-$\Lambda$CDM physics. The choice of modelling MG corrections to a pseudo cosmology, rather than a $\Lambda$CDM one, was outlined in Ref. Cataneo et al. (2019), which showed that the mass functions in the MG and pseudo cosmologies become quite similar, providing a smoother transition between inter- and intra-halo regimes. We direct the reader to Refs. Cataneo et al. (2020); Bose et al. (2020, 2022) for the construction of the reaction function using halo model quantities. Note that in the halo model reaction, 1-loop standard perturbation theory (SPT) is also employed to calibrate the 2-halo term to the 1-loop prediction, improving the overall prediction in the quasi-nonlinear regime. In this work we do not consider this calibration: Ref. Bose et al. (2022) showed that its effects are negligible for modifications to gravity inducing a scale-independent shift in the linear growth factor, as considered here.
@@ II.2 The nonlinear phenomenological parameterization
There are four additional parameters in Eqs. 10-12, characterising the nonlinear regime, $p_{1}$ $p_{4}$, representing MG phenomenological effects. The first one, $p_{1}$, parameterizes the strength of screening, $p_{2}$ and $p_{3}$ quantify the mass and environmental dependencies, respectively, whereas $p_{4}$ sets the scale of Yukawa suppression Yukawa (1935). The QNL correction, described by $S(k,a)$, was shown in Bose et al. (2022) to have a negligible effect for scale-independent MG theories, while it can have up to a 2% effect on the power spectrum for the Hu-Sawicki $f(R)$ model. We will ignore the QNL correction in the current implementation.
@@ IV Demonstrations
Parameter Flat Prior Fiducial
@@
First, we perform a test on mock data, generating synthetic DES-like galaxy-galaxy, galaxy-lensing and lensing-lensing (3$\times$2pt) correlations for a fiducial model with $\mu_{0}=0.4,\Sigma_{0}=0.1$ and $p_{1}=0.5$, and assuming a $\Lambda$CDM background expansion. The values of the other (cosmological) parameters are reported in Tab. 1, whereas the DES likelihood parameters are fixed at the standard values of DES Year-1 Abbott et al. (2019); Krause et al. (2017), to help us isolate the impact of including the nonlinear extension. As mentioned in Sec. III.1, the $p_{1}\rightarrow-\infty$ and $p_{1}\rightarrow\infty$ limits of the nonlinear parameter $p_{1}$ describe the fully screened (i.e. the power spectrum goes back to the $\Lambda$CDM prediction) and fully unscreened regimes, respectively. In practice, we have verified that varying $p_{1}$ over the prior range $p_{1}\in[-2,2]$ is sufficient to capture this variability , consistently with what shown in Ref. Tsedrik et al. (2024). This can be understood by noting that the transition from the screened to the unscreened regimes is modelled as an error function in the halo model reaction method Bose et al. (2020): at $p_{1}\pm 2$, the error function has largely reached its asymptotes.
The mock data are generated using MGCAMB. As our “data” covariance, we adopt the covariance matrix from the DES Year-1 likelihood, rescaled down by a factor of 25, as the actual DES Year-1 covariance would not allow us to constrain the nonlinear parameter $p_{1}$. We then run MCMC chains using MGCobaya, to see if we can recover the input model. To test our implementation, and to assess the added constraints brought in by the nonlinear scales, we run with and without the nonlinear extension. In the latter case, we only vary the linear parameters $\mu_{0}$ and $\Sigma_{0}$, whereas we include $p_{1}$ when the nonlinear extension is applied. In the analyses we run, we employ the priors reported in Tab. 1.
@@
To test the nonlinear extension to MGCAMB, we first check that it does not impact constraints derived from the linear scales alone. To do so, we analyse the DES-like data with and without the nonlinear extension, while applying the aggressive cut. We show the results of this test in Fig. 3 and the first two columns of Tab. 2: the recovered constraints on the parameters $\mu_{0}$ and $\Sigma_{0}$ are consistent with each other and with the fiducial model in both cases, demonstrating that indeed the nonlinear corrections does not impact the linear scales in unwanted ways. However, note that, when the nonlinear extension is disabled (green contours), the recovered values for the linear parameters $(\mu_{0},\Sigma_{0})$ are slightly shifted away from their true values: although small, this effect shows that nonzero values of the nonlinear parameter $p_{1}$ in the data might induce a bias in analyses that do not consider the effects of screening, even when limited to the linear scales.
@@
As a second test, we assess the improvement in the constraints when we include the nonlinear scales. For this, we analyse the baseline DES-like synthetic data with and without the nonlinear extension. We also consider the case where $p_{1}$ is held fixed at the fiducial value. The joint constraints on $\mu_{0}$, $\Sigma_{0}$ and $p_{1}$ are shown in Fig. 4. It is clear that we obtain better constraints on $\mu_{0}$ and $\Sigma_{0}$ when the nonlinear extension is used, even after marginalizing over $p_{1}$. Since $p_{1}$ and $\mu_{0}$ do not degenerate strongly, the constraint on $\mu_{0}$ is only slightly improved by fixing $p_{1}$. On the other hand, the constraint on $\Sigma_{0}$ becomes tighter by fixing $p_{1}$ due to the breaking of the degeneracy between $\Sigma_{0}$ and $p_{1}$. This demonstrates that the improvement is coming largely from the added scales in the “DES baseline” dataset compared to “aggressive cut”. We also note that there is a bias in the recovered value of $\Sigma_{0}$ when the linear cut is applied. We report the parameter values we recover in Tab. 2, showing that the $1\sigma$ uncertainty on $\mu_{0}$ is reduced by a factor of $\sim 2$ when the nonlinear extension is used, which agrees with Fig.3. of Ref. Tsedrik et al. (2024).
Figure 4: Comparison of the constraints on $\{\mu_{0},\Sigma_{0}\}$ obtained with and without the nonlinear extension. In the latter case, the aggressive cut is applied. The constraints are derived from synthetic data generated from the fiducial parameters reported in Tab. 1. Constraints are tighter when the nonlinear extension and the baseline data are used. We also show the constraint obtained by fixing $p_{1}$ to its fiducial value with the DES baseline synthetic data.
Parameter Flat Prior Posterior
@@
$\mu_{0}$ (-1.0, 1.0) 0.003 $\pm$ 0.201 -0.012 $\pm$ 0.216
$\Sigma_{0}$ (-0.5, 0.5) 0.016 $\pm$ 0.039 0.043 $\pm$ 0.045
$p_{1}$ (-2, 2) -0.0006 $\pm$ 1.1283 -
$\Omega_{b}h^{2}$ (0.005, 0.1) 0.02245$\pm$ 0.00014 0.02247$\pm$ 0.00014
$\Omega_{c}h^{2}$ (0.001, 0.99) 0.1190$\pm$ 0.0009 0.1187$\pm$ 0.0009
$n_{s}$ (0.8, 1.2) 0.9692$\pm$ 0.0036 0.9690$\pm$ 0.0037
$\mathrm{ln}(10^{10}A_{s})$ (1.61, 3.91) 3.048$\pm$ 0.015 3.042$\pm$ 0.015
$100\,\theta_{\mathrm{MC}}$ (0.5, 10) 1.0410$\pm$ 0.0003 1.0411$\pm$ 0.0003
$\tau$ (0.01, 0.8) 0.057$\pm$ 0.008 0.055$\pm$ 0.008
Table 3: The priors and the mean parameter values with the $68\%$ CL uncertainties derived from the combination of Planck 2018, BAO, RSD, and SN Ia (PBRS) data with DES Year-3 baseline data, with and without the nonlinear extension (also see Fig. 5). When excluding the nonlinear extension, we use the same linear cut as the DES collaboration Abbott et al. (2023). Note the improvement on the constraints on the MG parameters when the nonlinear extension is included.
@@
Having tested our setup, we apply the extended MGCAMB to current cosmological datasets, deriving constraints on $\mu_{0}$ and $\Sigma_{0}$, along with the main cosmological parameters. For this, we use the DES Year-3 3$\times$2pt data Abbott et al. (2023), Planck 2018 CMB temperature, polarization and lensing power spectra Aghanim et al. (2019), joint measurements of baryon acoustic oscillations (BAO) and redshift-space distortions (RSD) from eBOSS DR16, the SDSS DR7 MGS data Ross et al. (2015), the BAO measurement from 6dF Beutler et al. (2011), and the Pantheon sample of uncalibrated supernovae Scolnic et al. (2018). In addition to $\mu_{0}$ and $\Sigma_{0}$, we vary all the standard cosmological parameters: the baryon density $\Omega_{b}h^{2}$, the CDM density $\Omega_{c}h^{2}$, the amplitude $A_{s}$ and the spectral index $n_{s}$ of the primordial fluctuations spectrum, the angular size of the sound horizon at decoupling $\theta_{\mathrm{MC}}$ and the optical depth $\tau$. To assess, the impact of including nonlinear scales, we perform two analyses, with and without the nonlinear extension, applying the DES Year-3 linear cut in the latter. We show the recovered posteriors for $\mu_{0}$, $\Sigma_{0}$ and $p_{1}$ in Fig. 5: it is clear that we obtain stronger constraints on $\mu_{0}$ and $\Sigma_{0}$ with the nonlinear extension. In fact, whilst the posterior on $\mu_{0}$ is largely unaffected, the improvement on the $\Sigma_{0}$ constraint is apparent. We can understand this by noting that DES is primarily a weak lensing survey, which primarily constraints $\Sigma$. The priors and the mean values with uncertainties for all parameters are provided in Tab. 3. The constraints we obtain on $\mu_{0}$ and $\Sigma_{0}$ with the DES Year-3 linear cut are consistent with those reported by the DES collaboration Abbott et al. (2023).
Finally we make a note on baryonic effects, known to play a large role in the nonlinear regime. These effects have been found to be largely degenerate with the modified-gravity screening parameter $p_{1}$ in Ref. Tsedrik et al. (2024), where the authors provided forecasts much deeper into the nonlinear regime than considered here. As this work is meant to be a demonstration of the new implementation in MGCAMB, we do not consider very nonlinear scales and we do not include baryonic effects. We note that the baseline scale cut developed by the DES collaboration was chosen to avoid contamination from baryonic effects. This can be confirmed by the fact that $p_{1}$ is not well constrained by the DES Year-3 3x2pt data. However, with the covariance rescaled by a factor of 25, $p_{1}$ is well constrained, and this means that baryonic effects will play a significant role. It should be noted that it may be sufficient to include a $\Lambda$CDM-derived baryonic boost factor along with the prediction of $P_{\rm NL}(k,z)$ even in MG scenarios, as supported by various recent studies (Mead et al., 2016; Arnold and Li, 2019; Hernández-Aguayo et al., 2021), which show that baryonic and MG effects are largely decoupled.